OpenStudy (anonymous):

ANSWER THIS PLEASE 7d + 12 – 4d – 3 (2 points) 19d – 7 3d + 9 3d^2 + 9 12d

OpenStudy (anonymous):

7d-4d=3-12 3d=-9 d=-9/3 d=-3

OpenStudy (anonymous):

wait so what answer?